### v1

@@ 1 Introduction
A clinician’s daily workflow is fundamentally an exercise in resource allocation, split between high-value complex clinical judgment and repetitive administrative tasks involving the retrieval and synthesis of fragmented data from the Electronic Health Record (EHR). This clerical burden creates a significant operational bottleneck; it is a primary driver of physician burnout, which has been linked to significant patient safety risks and estimated costs to the U.S. healthcare system reaching billions of dollars annually (goh2019; sinsky2022burnoutcost; muir2022nurseburnout; tawfik2024burnout). Integrating technology to optimize these workflows and alleviate cognitive load has shown promise in improving clinician well-being, operational efficiency, and clinical outcomes. In particular, the advent of Large Language Models (LLMs) has opened new avenues for such enhancement, ranging from ambient AI scribes that reduce documentation time (tierney2024ambient; tierney2025ambient) to “copilots” that verify medication prescriptions (medic-mohsen) or selectively trigger safety alerts in live care (korom2025penda).
In this paper, we focus on one such high-frequency administrative task: the calculation of standardized medical risk and stratification scores. These calculations are ubiquitous in hospital settings, for instance, determining pneumonia severity (CURB-65) upon a patient’s arrival at an Emergency Department, or assessing stroke risk (CHA2DS2-VASc) to guide anticoagulation therapy (lim2003curb65; lip2010cha2ds2vasc). These metrics synthesize patient history, vitals, and laboratory results into actionable scalars that directly inform high-stakes operational decisions: prioritizing patient flow in triage, determining the appropriate level of care (e.g., admission planning), and applying preventative interventions (prophylaxis).
Traditionally, these scores are computed manually: a clinician reads the EHR, extracts the necessary criteria, and inputs them into a bedside calculator such as MDCalc, a web app that has between three and five million visits per month and is used by 65% of physicians in the US (mdcalc). However, the manual extraction of clinical criteria from unstructured notes is time-consuming and prone to mistakes. The emergence of LLMs offers a transformative vision: automated systems that can read EHR notes, extract relevant clinical features, and reliably compute these scores end to end (xie2024preliminary; goodell2025; agentmd; wang2025agent; roeschl2025). But as we move toward deploying or even researching such systems, a fundamental question of quality assurance arises: how do we ensure these automated systems are computing the “right” scores, specifically, scores that align with the ground truth judgment of a licensed physician?
To answer this question, the research community requires an accessible evaluation standard. In this domain, MedCalc-Bench (medcalc) stands as a foundational contribution. By curating de-identified patient notes from PubMed Central and mapping them to 55 widely calculated medical scores on MDCalc web app, medcalc established the first large-scale public testbed, involving 11,000 samples, for end-to-end clinical score computation. To generate labels at scale, the authors employed a two-stage pipeline: using GPT-4 to extract clinical features from text, followed by Python scripts to aggregate those features into final scores. This effort provided the field with a standardized definition of the task and a shared dataset, enabling subsequent research to benchmark progress in medical reasoning (oh2025; riskagent). Consequently, the labels provided by MedCalc-Bench have effectively become the ground truth against which new models are evaluated and, increasingly, the target distribution on which they are trained.
However, the reliance on largely LLM-generated “gold standard” labels introduces a unique systemic risk: these labels inherently echo the model’s capabilities and flaws at the time of creation. The original MedCalc-Bench labels relied on the extraction capabilities of GPT-4 in early 2024 and the assumption that programmed calculators perfectly mirror clinical intent. When such a benchmark is treated as a static oracle, errors in the original extraction or calculator logic become enshrined as systemic biases in the evaluation signal. This creates a dangerous feedback loop, particularly when benchmarks transition from evaluation sets to training signals. In modern Reinforcement Learning (RL) setups (ahmadian2024), inaccurate labels are not just a mislabeled yardstick, but a biased teacher that actively steers models away from clinical reality.
While benchmarks have been instrumental in accelerating progress, this challenge of latent inaccuracies necessitates a more rigorous approach to their lifecycle, grounded in principles of continuous quality control. We argue that LLM-assisted benchmarks for complex clinical workflows are not static measurement instruments; they are socio-technical artifacts whose labels inherit the failure modes of the models, tools, and procedural assumptions used at the time of creation. They should therefore be treated as living documents with explicit maintenance practices (auditing, clinician adjudication, versioning, and documentation). We do not position our improved labels as a definitive replacement for MedCalc-Bench; rather, we use the original release as a case study to motivate transparent, repeatable benchmark stewardship and to quantify how label fragility affects downstream model development.
In this work, we approach benchmark maintenance not as an ad-hoc error correction task, but as an optimized quality assurance process. We design a systematic, replicable collaboration between advanced agentic LLMs (as verifiers) and human physicians, focused on the central operational constraint: scarcity of expert attention. Our approach proceeds in three steps: (1) an automated audit where an LLM verifier flags potential errors in the original labels; (2) an independent maintenance pipeline that recomputes scores from scratch to identify systematic divergences; and (3) a targeted physician adjudication of divergent cases. In short, we leverage agentic AI to scale re-auditing and triage, thereby concentrating scarce clinician time on the most consequential disagreements.
The above approach shows that a non-trivial fraction of the original labels diverge from physician judgment due to feature extraction errors, calculator logic mismatches, or clinical ambiguity (e.g., unanswerable questions). Furthermore, we demonstrate the downstream impact of unmaintained labels through a controlled LLM fine-tuning experiment. By fine-tuning a Qwen3-8B model (qwen3) via Group Relative Policy Optimization (GRPO) (grpo), a commonly used RL algorithm, we show that changing only the reward labels (original versus maintained) yields an 8.7% absolute difference in test accuracy when evaluated against physician-aligned labels. The goal of this experiment is not to claim superior benchmark performance; rather, it is to test whether label noise materially impacts trained model accuracy, a meaningful question because one plausible outcome was that the noise would be “washed out” by the strength of signal in the data. Our results indicate otherwise, confirming that benchmark maintenance has tangible consequences for downstream alignment.
Our contributions are summarized as follows:
  • Modeling the label generation process. We mathematically model the data generating process underlying the original MedCalc-Bench gold labels (§2.1). This abstraction allows us to attribute how procedural errors in extraction and aggregation distort the ground truth estimate. Furthermore, we extend this model to the alignment phase, formalizing how label noise propagates into the reward function of the Markov Decision Process (MDP) used in RL training (§4.1).
  • Operationalizing benchmark stewardship. We design and implement a physician-in-the-loop system that operationalizes scalable oversight by optimizing the allocation of expert attention. By aggregating evidence from two independent, agentic LLM pipelines, the system triages the most contentious instances in MedCalc-Bench for targeted physician expert review (§3.1, 3.2). In a single-blind validation sample, the resulting maintained labels exhibit higher agreement with physician judgment (§3.2.1) [^1] Most raw results and intermediate data are published in https://github.com/junzeye/validate-medcalc-labels; we will update the repository with processed code and data artifacts as curation and formatting is completed..
  • Quantifying the impact on LLM alignment. Using a controlled RL experiment on a modern open-source LLM, we quantify how different label choices (original versus maintained) become training rewards that significantly impact LLM alignment (§4). This analysis highlights that label fragility in an LLM-assisted benchmark can simultaneously distort model behavior and the metrics by which that behavior is perceived by its human trainers.
1.1 Literature Review
Integrating LLMs into clinical workflows.
[…]

### v2

@@ 1 Introduction
A clinician’s day often alternates between patient care, time spent on complex judgment, and time spent on repetitive “information work”: retrieving scattered facts from the electronic health record (EHR), reconciling fragmented documentation, and translating that information into standardized actions. This clerical burden contributes to burnout and has been linked to patient-safety and health-system consequences.[14, 37, 27, 39] Tools based on Large Language Models (LLM) are increasingly proposed to reduce this burden, including ambient documentation support and targeted safety guardrails in healthcare delivery.[40, 1, 30, 18, 8]
Routine computation of medical risk and severity scores is another common information processing scenario that appears suitable for automation. Risk scores such as LACE, CURB-65, and CHA2DS2-VASc condense patient history, vital signs, and laboratory results into numeric values that inform triage, admission planning, and preventive therapy;[42, 20, 22] a worked out example in Supplementary Appendix §D.1 illustrates a computation routine. In practice, clinicians extract criteria from narrative notes or structured EHR data, and then manually calculate the final scores or enter information into calculators such as MDCalc.[26] Manual extraction is time-consuming and error-prone, motivating the adoption of large language model (LLM) systems that can read notes and compute scores end-to-end.
As these automated systems are developed, evaluating their efficacy in a scalable but also clinically aligned manner becomes a safety challenge: how do we responsibly verify that they compute scores as accurately as would a conscientious human physician? Public task benchmarks are an important first step for transparency and reproducibility, but large-scale benchmark construction is increasingly assisted by LLMs themselves through feature extraction and normalization. As a prime example, MedCalc-Bench [16] is a widely used large-scale public benchmark for end-to-end medical risk score computation. Besides popularity in academic research,[6, 23, 21, 47, 43, 28] it was one of two medical benchmarks evaluated on by Anthropic in its recent launch of the “Claude for Healthcare” initiative.[3, 4]
MedCalc-Bench was constructed from de-identified PubMed Central case reports paired with 55 commonly-used medical calculators on MDCalc. Its original labeling pipeline used GPT-4 for feature extraction and scripted aggregation for final score computation. This automated labeling pipeline yielded a large task dataset of 11,000 labeled instances. The scale and comprehensive calculator coverage of MedCalc-Bench make it an important contribution to clinical natural language processing (NLP) research, as AI model training is inherently data-hungry. Manually producing benchmark labels at this scale is difficult, as sample-by-sample clinician adjudication is slower and much more expensive than LLM API calls: sizes of fully physician-labeled medical score datasets, such as [11] and [32], are in the scale of low hundreds. However, the expediency introduces a structural vulnerability: labels can inherit systemic failure modes of the language models, tools, and procedural assumptions used to produce reference labels. It is a concerning situation because in effect, these synthesized benchmark labels are not only the “ground truth” used to measure progress and, increasingly, also a target distribution for AI developers to align their models towards, for example via reinforcement learning (RL).[29, 2] When the same labels are reused as RL reward signals, label errors are no longer only measurement noise: they can actively shape the behavior of future models, turning the benchmark from a miscalibrated yardstick into a biased teacher.
In this paper, we audit and improve the clinical soundness of reference labels originally released by MedCalc-Bench in its NeurIPS 2024 publication.[16] Schematically, we propose a scalable stewardship protocol in which agentic LLM pipelines—LLM-based verifiers instructed to follow a structured checklist and use tools such as search and computation—audit existing labels and independently recompute scores; automated triage then concentrates scarce physician time on the highest-impact disagreements. Physician review resolves contention and identifies truly unanswerable or mismatched note-question pairs, for which abstention (e.g., “unknown”) is safer than forced computation. Importantly, our goal is not to fault the creators of MedCalc-Bench; rather, we use their original dataset release as a motivating case study to develop a transparent, repeatable stewardship process for benchmarks that increasingly influence both evaluation and post-training. Our contributions are as follows:
  • Propose systematic stewardship for clinical workflow benchmarks as safety infrastructure and describe the risks of treating partially LLM-generated labels as static oracles.
  • Introduce a physician-in-the-loop protocol that combines agentic audit, independent recomputation, and automated triage, with clinician adjudication for disputed cases.
  • Apply the protocol to MedCalc-Bench, characterize major sources of label failure, and incorporate an abstention label for unanswerable note-question pairs.
  • Demonstrate in a controlled RL setting that correcting for latent training label bias can materially change LLM behavior and performance interpretation.
The remainder of the manuscript describes the benchmark’s original label-generation process and label failure modes, presents our stewardship protocol and physician review process, and then examines how revised labels change evaluation and post-training conclusions.
2 Methods
2.1 Benchmark setting and task definition
MedCalc-Bench is a public benchmark designed to evaluate whether language models can compute common medical scores from free-text narratives, without requiring external calculator use.[16] Each benchmark instance presents a model with a clinical context, typically several paragraphs of a de-identified patient narrative, and a question requesting a specific medical score (e.g., “What is the patient’s HAS-BLED score?”). The model must return either a numeric value, a datetime, or, in our extended formulation, an explicit abstention ($\mathsf{N/A}$”) when the context lacks necessary information or is clinically mismatched to the question. The benchmark contains 11,000 context-question pairs (10,053 training; 1,047 test) covering a set of 55 different score questions. This task mirrors what clinicians do: read documentation, identify the criteria a particular score requires, and apply the scoring rule. The benchmark therefore tests whether a model can perform both subtasks, criterion extraction from unstructured text and rule-based aggregation, correctly and in sequence.
Dataset versioning.
Our analysis focuses on the peer-reviewed release of MedCalc-Bench (v1.0), which serves as the standard baseline for current academic and industry evaluations (e.g., [6], [3]; see also the last paragraph of Appendix §D.2). We note that while the authors began releasing updates (v1.1, v1.2) during the final preparation of this manuscript, our auditing and reinforcement learning experiments were completed prior to these releases. We therefore utilize v1.0 to ensure: (1) comparability with the existing body of concurrent work, and (2) that our stewardship protocol is tested against the distribution of errors present in the widely circulated version of the benchmark. A detailed analysis of the performance discrepancies between dataset versions is provided in Supplementary Appendices B and G.
Label generation and correctness criteria.
In MedCalc-Bench’s publication[16], reference labels were produced by a three-stage procedure: GPT-3.5 matched patient contexts to appropriate calculator questions, GPT-4 extracted clinical features from each narrative (e.g., patient demographics, symptoms, laboratory values), and custom Python scripts aggregated those features according to each calculator’s formula or decision logic. At test time, a model prediction is scored correct if it falls within $\pm$5% of the reference label for numeric outputs or matches exactly for datetimes. We extend this scheme to treat abstention as correct when our recomputed reference label is itself “unknown.”
Reference label as an estimator.
We define physician-aligned ground truth, $y^{\star}$, as the answer a conscientious physician would provide for a given note-question instance $(C,q)$. The corresponding benchmark reference label, which we denote by $\hat{y}$, is an estimate of this (latent) ground truth; $\hat{y}^{\mathrm{original}}$ is the reference label originally given by MedCalc-Bench, while $\hat{y}^{\mathrm{new}}$ is the label our agentic workflow independently recomputes in §2.4. Estimation error can enter at any of the three stages, whether through inappropriate context-calculator pairing, hallucinated or misread feature values, or aggregation scripts that diverge from published scoring rules. Task-level ambiguity, such as missing timepoints or guideline version differences, introduces additional uncertainty. Recognizing these labels as fallible estimates, rather than one-off oracles, motivates the stewardship process described below.
Notation for label indexing.
Throughout this paper, $y$ denotes a single label for an instance $(C,q)$, regardless of whether it contains a subscripted index. When discussing multiple instances, we use subscripts ($y^{\star}_{i}$, $\hat{y}^{\mathrm{original}}_{i}$, or $\hat{y}^{\mathrm{new}}_{i}$) to disambiguate the label associated with instance $i$. When the instance is clear from context (e.g., there is only one $(C,q)$), we drop the “i” subscript and write $y^{\star}$, $\hat{y}^{\mathrm{original}}$, or $\hat{y}^{\mathrm{new}}$ for readability.
Figure 1: Our benchmark stewardship utilizes two distinct LLM agent workflows (Phases 1 & 2) to assure MedCalc-Bench’s label quality. Besides their prompt and output type (Yes/No verdict versus a recomputed label), the two workflows differ in what the LLM agent is given as context: the Phase 1 agent is shown the original reference label and its derivation metadata, whereas the Phase 2 agent is only provided the patient narrative and score question; the latter omission is intended to reduce any anchoring bias towards $\hat{y}^{\mathrm{original}}$. In Phase 3, three physician authors (DT, AG, NK) independently compute 50 instances where $\hat{y}^{\mathrm{original}}$ and $\hat{y}^{\mathrm{new}}$ show high disagreement. sMAPE stands for symmetric Mean Absolute Percentage Error (Eq. 2); it measures the relative disagreement in $\{(y^{\star}_{i},\hat{y}^{\mathrm{new}}_{i})\}_{i=1}^{50}$ and $\{(y^{\star}_{i},\hat{y}^{\mathrm{original}}_{i})\}_{i=1}^{50}$.
2.2 Stewardship process overview
The central constraint in validating benchmark reference labels is not machine computation but physician experts’ attention. GPU hours are elastic; physician hours are not. Our process is therefore designed around a principle of phased resource commitment: each phase produces evidence that justifies, or gates, incremental investment, concentrating scarce expert time on instances most likely to benefit from it. This design balances scalability through LLM-based meta-evaluation with the authority of physician judgment, which remains the ultimate source of clinical ground truth but is routed to high-information decisions.
The process proceeds in three phases, in which the first two serve as an automated instances triage. Phase 1 (automated audit) applies a tool-augmented LLM verifier to existing labels, flagging instances where the original derivation of $\hat{y}^{\mathrm{original}}$ appears clinically unsound. This phase is inexpensive to run and does not commit to replacement labels; its purpose is to provide an early signal of the overall error rate, justifying further investigation. Phase 2 (independent recomputation) deploys a separate LLM pipeline that sees only the patient context and question, not the original label or its provenance, and recomputes a label $\hat{y}^{\mathrm{new}}$ from first principles. Disagreement between the recomputed label and the original, beyond a clinically meaningful threshold, identifies cases where at least one estimate is wrong. Finally, Phase 3 (physician adjudication) directs clinician review to the subset of instances where Phase 2 recomputation diverges most substantially from original labels, maximizing information gained per expert hour.
2.3 Phase 1: Tool-augmented automated audit of original labels
Objective.
Phase 1 is a low-cost automated audit of MedCalc-Bench test labels. Its purpose is to provide an early signal of whether the benchmark’s original label-derivation procedure produces clinically meaningful inconsistencies, and hence whether it is worth investing in the more resource-intensive Phases 2–3. The fraction of $\{\hat{y}^{\mathrm{original}}_{i}\}_{i=1}^{1047}$ identified in this phase as problematic also serves as a sensitivity check of the problematic fraction implied by independent label recomputation in Phase 2, which has a different agentic workflow setup.
Audit inputs and label-derivation metadata.
We built an LLM auditor agent for the MedCalc-Bench test split (1,047 instances). For each instance, the auditor received: (1) the patient context $C$ (a de-identified clinical narrative), (2) the calculator question $q$, (3) the original benchmark label $\hat{y}^{\mathrm{original}}$, and (4) original metadata describing how MedCalc-Bench derived $\hat{y}^{\mathrm{original}}$. (4) includes the structured clinical features extracted from $C$ by the original pipeline (e.g., laboratory values, vitals, demographics) and a description of the calculator-specific aggregation logic used to map extracted features to $\hat{y}^{\mathrm{original}}$. Supplying these artifacts enables the auditor to evaluate whether $\hat{y}^{\mathrm{original}}$ is consistent with both the clinical narrative and a clinically sound extraction-and-aggregation procedure for score question $q$.
Auditing agent and tool access.
We implemented the auditor as a reasoning language model (Gemini 2.5 Pro, June 2025 version) configured as a verifier with web-search access. During each audit, the agent could retrieve calculator definitions, rule tables, and guideline statements as needed, improving factual grounding beyond parametric knowledge.
Audit procedure and conservative flagging.
Each Phase 1 audit run was designed to verify the benchmark’s existing label and its derivation artifacts, not to generate a new blinded label. Concretely, the auditor (i) identified the appropriate scoring rule for $q$, using web search when needed; (ii) checked whether the required clinical criteria are present and interpretable in $C$; (iii) validated that the structured feature values reported in the benchmark metadata agree with the evidence stated in $C$; and (iv) performed a sanity-check calculation under the identified rule. The agent was instructed to be conservative and to flag Suspicious only when it could articulate a clinically meaningful inconsistency.
Reliability controls via repeated sampling and supermajority voting.
We use a sampling temperature of 1 for generating outputs, following Google’s official recommendation.[12] Because LLM outputs are stochastic, we ran the auditing agent five times independently for each $(C,q,\hat{y}^{\mathrm{original}})$ sample using identical instructions. We then aggregated verdicts with a conservative supermajority rule: an instance was labeled Likely Error only if at least 4 of 5 runs flagged $\hat{y}^{\mathrm{original}}$ as incorrect in a clinically significant way; otherwise it was labeled Passed. This design is intended to prioritize precision over recall.
Clinician spot-check.
To assess if the automated audit’s flags were clinically sensible, a physician author (AG) spot-checked seven cases flagged as Likely Error within his clinical domain. For each selected case, the physician reviewed the patient context, the calculator question, the original benchmark label, and the auditor’s written critique, and recorded whether the critique reflected a clinically meaningful label problem.
2.4 Phase 2: Independent recomputation of reference labels
Objective and Scope.
Phase 2 proposes revised reference labels by independently recomputing each score from first principles, without access to the original benchmark label or its derivation artifacts. This phase aims to (i) construct a set of high-confidence updated labels for evaluation and downstream use, and (ii) quantify disagreement between independently recomputed labels and the original labels to triage cases for clinician adjudication in Phase 3. We applied the Phase 2 relabeling pipeline to the full MedCalc-Bench test split (n=1,047n=1{,}047) and to a stratified random sample of the training split (n=5,183n=5{,}183, ${\sim}50\%$ of the original 10053-instance training set) stratified by question type $q$.
Independent labeling agent with comprehensive tools access.
We implemented a separate relabeling agent (Gemini 2.5 Pro) configured as a tool-augmented clinical calculator. For each instance, the agent received only the patient context and calculator question $(C,q)$ (not seeing $\hat{y}^{\mathrm{original}}$ nor any released extraction/aggregation metadata), to reduce anchoring to the original pipeline. The agent had access to (i) a Google Search API to retrieve authoritative calculator definitions and scoring rules (e.g., MDCalc and guideline documents), and (ii) a Python code execution sandbox for arithmetic and unit conversions.
The agent was instructed to (1) identify the appropriate scoring rule, (2) extract all relevant criteria from $C$, (3) map extracted evidence to score components, and (4) compute the final value. When required inputs were missing or the question could not be resolved clinically from the provided context, the agent was instructed to abstain ($\mathsf{N/A}$”) rather than infer or hallucinate. For reliable parsing, the agent was required to enclose its final output in a strict regex-readable format (a scalar numeric value, a datetime, or $\mathsf{N/A}$”). The complete system and user prompts are provided in Supplementary Appendix §A.
Supermajority-Based Labeling.
To ensure stability and reliability, each instance is processed by five independent copies of the Gemini agent at a temperature of 1. We employ a supermajority voting rule: a “high-confidence” label ($\hat{y}^{\mathrm{new}}$) is recorded if at least four out of the five independent runs agree on the output type and value. [^1] To balance stability with data retention, we also retain instances with a 3/5 supermajority if at least one of the remaining two outputs falls within a proximity threshold: $\pm 5\%$ for numeric values or $\pm 1$ day for dates. This “near-consensus” logic reduces unnecessary deferrals from minor variations in agent outputs. Instances failing to reach this quorum are marked as “Deferred” and excluded from the high-confidence set. For real-valued numeric outputs, agreement for the majority is defined after rounding to two decimal places; for datetimes and categorical outputs, an exact match is required.
Automated Triage for Physician Review.
To identify high-impact label failures for targeted clinician recomputation in Phase 3, we compare $\hat{y}^{\mathrm{new}}$ against $\hat{y}^{\mathrm{original}}$ for each instance. We define a symmetric discrepancy metric, rel.err, for the two label estimators:
$\textsf{rel.err}(\hat{y}^{\mathrm{original}},\hat{y}^{\mathrm{new}}):=\frac{|\hat{y}^{\mathrm{original}}-\hat{y}^{\mathrm{new}}|}{\max(|\hat{y}^{\mathrm{original}}|,|\hat{y}^{\mathrm{new}}|)}$ (1)
Following the $\pm 5\%$ tolerance used by MedCalc-Bench for grading model outputs, an instance is flagged as a “Likely Error” if $\textsf{rel.err}(\hat{y}^{\mathrm{original}},\hat{y}^{\mathrm{new}})>0.05$. Cases where either label is N/A or datetime are handled respectively by exact match and specific time-threshold differences.
2.5 Phase 3: Targeted physician adjudication
Objective.
Phase 3 concentrates independent physician recomputation and review efforts on the subset of test instances with the largest disagreement between the original label ($\hat{y}^{\mathrm{original}}$) and the Phase 2 recomputed label ($\hat{y}^{\mathrm{new}}$), thus providing a targeted adjudication.
Triage and selection.
Physician authors (DT, AG, NK) first screened calculator questions $q$ for respective domain fit. Author JY ranked test instances by disagreement between $\hat{y}^{\mathrm{original}}$ and $\hat{y}^{\mathrm{new}}$, treating any abstention ($\mathsf{N/A}$) by either label as maximal disagreement, and then sampled 50 instances from the ranked list, assigning each instance to a physician aligned with the score type. Each selected instance was reviewed once by a single physician.
Single-blind physician recomputation.
For each assigned instance, the physician received only the patient context and question $(C,q)$ and independently recomputed the score from first principles in most of these, without seeing $\hat{y}^{\mathrm{original}}$ or $\hat{y}^{\mathrm{new}}$ (single-blind to reduce anchoring). Physicians could consult external references (e.g., MDCalc or guideline documents) and could abstain ($\mathsf{N/A}$) when the score was not clinically computable from the context. Physicians optionally recorded brief free-text comments (e.g., missing inputs, ambiguity, guideline/version assumptions). We denote a physician-computed label by $y^{\star}$.
Agreement criteria, error metrics, and uncertainty.
We compare $\hat{y}^{\mathrm{original}}$ and $\hat{y}^{\mathrm{new}}$ to $y^{\star}$ using (i) an agreement indicator and (ii) symmetric mean absolute percentage error (sMAPE, [5]) for answerable numeric cases. For real-valued outputs, agreement is defined as relative deviation within $\pm 5\%$ (i.e., $|\hat{y}-y^{\star}|/|y^{\star}|\leq 0.05$). For integer-valued ordinal scores, agreement is defined as $|\hat{y}-y^{\star}|\leq 1$ when $y^{\star}<20$. For abstentions, agreement requires an exact match (both $y^{\star}$ and $\hat{y}$ equal $\mathsf{N/A}$). For $N$ instances, sMAPE is computed as:
$\mathrm{sMAPE}(\{y^{\star}_{i}\}_{i=1}^{N},\{\hat{y}_{i}\}_{i=1}^{N})=\frac{100\%}{N}\sum_{i=1}^{N}\frac{2|y^{\star}_{i}-\hat{y}_{i}|}{|y^{\star}_{i}|+|\hat{y}_{i}|}.$ (2)
We estimate 95% confidence intervals for sMAPE via bootstrapping 10,000 resamples.
Post-adjudication label finalization and qualitative analysis.
We used physician recomputations to finalize labels on the reviewed subset, yielding $\hat{y}^{\mathrm{final}}$ where $\hat{y}^{\mathrm{final}}=y^{\star}$ for physician-reviewed instances and $\hat{y}^{\mathrm{final}}=\hat{y}^{\mathrm{new}}$ otherwise. For instances where $\hat{y}^{\mathrm{final}}\neq y^{\star}$, we examine reasoning traces of the Gemini labeling agent, and if a systematic bias is identified relative to the physician’s ground truth (e.g. scoring criteria of some $q_{i}$), we analogously revise $\hat{y}^{\mathrm{new}}$ labels for unreviewed $(C,q_{i})$ instances. We also perform a post-hoc qualitative coding of physicians’ free-text comments to summarize recurrent causes of disagreement. We release both $\hat{y}^{\mathrm{new}}$ (pre-adjudication) and $\hat{y}^{\mathrm{final}}$ (post-adjudication) with triage metadata.
2.6 Controlled Study of Label Quality’s Downstream Impact
To determine if reference label inaccuracies are just benign “white noise” or a directional distortion that misaligns model behavior, we performed a controlled reinforcement learning (RL) experiment. This simulates the potential consequence when an AI researcher optimizes for biased reference labels that reflect synthetic artifacts rather than clinical reality.
We finetuned two identical copies of Qwen3-8B, an open-weights LLM with 4.6 million downloads from Hugging Face in August 2025,[31] using Group Relative Policy Optimization (GRPO) which is a modern RL algorithm widely used in LLM alignment.[34] The algorithm iteratively modifies a language model’s weights to increase the probability of generating completions that receive a higher reward. In this framework, the reference label ($y^{\star}$, surrogated in practice by $\hat{y}$) directly specifies a rule-based reward function; changing $\hat{y}$ is therefore functionally equivalent to changing the model’s implicit alignment objective.
To isolate the causal effect of reference label choice, all training factors, including the base model, prompt templates, hyperparameters, and compute budget, were held constant while varying only the source of the training reward:
  • Control Arm (Original Labels): One copy of Qwen3-8B was trained to maximize agreement with the original MedCalc-Bench reference labels ($\hat{y}^{\mathrm{original}}$).
  • Treatment Arm (Updated Labels): Another copy of Qwen3-8B was trained to maximize agreement with our Phase 2 independently recomputed labels ($\hat{y}^{\mathrm{new}}$).
More technical details are given in Supplementary Appendix §D.4 and §E.
Both models were trained on 4,593 training instances and evaluated on a held-out test set of 887 instances. To test the sensitivity of evaluation outcomes to the choice of reference labels, the primary outcome measure was the accuracy of both models’ computed scores when graded against the updated test labels ($\hat{y}^{\mathrm{new}}$).
Data and Code Availability
We have posted the raw outputs of Phases 1 and 2 workflows, high-confidence recomputed labels of MedCalc test set instances and train set instances obtained in Phase 2, physician adjudication results in Phase 3, and code to reproduce the RL experiment. To access, visit the GitHub repository: https://github.com/junzeye/validate-medcalc-labels.
3 Results
3.1 Phase 1
Audit statistics of §2.3.
Running five independent copies of the auditor agent on MedCalc’s test set (n=1,047) surfaced a substantial number of suspicious labels: out of 1,047 instances, the $\hat{y}^{\mathrm{original}}$’s of 279 (26.6%) were flagged as Likely Error by the audit, i.e. marked as Suspicious by at least 4 out of 5 auditor reviews. Flagged issues were distributed broadly across the different questions $q$, covering 40 out of the total 55 distinct scoring questions in the test set.
Figure 2: Representative error types. (a) Feature extraction error: GPT-4 might have confused “hemoglobin” with “albumin”, extracting a value that is physiologically impossible; (b) Incorrect aggregation logic: an incorrect Python code for Glasgow Coma Scale aggregation that double-counts a feature value, inflating $\hat{y}^{\mathrm{original}}$; (c) $q$ is not answerable given $C$: a Sodium correction for hyperglycemia inappropriately applied to a hypoglycemic patient.
Qualitative error taxonomy.
Manual inspection of flagged instances reveals three recurring failure modes in the original ground truth derivation: (a) feature extraction errors, (b) incorrect aggregation logic, i.e. flawed scoring rules, and (c) $q$ is not answerable given $C$. Figure 2 illustrates one example of each. In the case of (c), it would be more reasonable to reward the LLM for abstaining from answering the impossible and responding with $\mathsf{N/A}$ (“Not Available”), instead of hallucinating an answer. Moreover, instead of throwing away all these instances, we suggest that the benchmark recycle them by keeping $(C,q)$ but revising the ground truth label from $\hat{y}^{\mathrm{original}}$ to $\mathsf{N/A}$ (see Supplementary Appendix §F for two examples). Correspondingly, the benchmark’s inference prompt could acknowledge that there are instances where the correct answer is to abstain. A safely useful LLM system, we believe, should act as a triage agent that seeks human review for out-of-distribution $(C,q)$ instances, while strictly limiting abstention on normal cases to preserve the efficiency gains of automation.
Clinician spot-check feedback.
A physician author (AG) independently selected and reviewed 7 Suspicious instances related to his clinical specialty. For each instance, the physician read both $(C,q,\hat{y}^{\mathrm{original}})$ and the auditor agents’ critique of $\hat{y}^{\mathrm{original}}$. They agreed with the error identification in 7/7 reviews. The sample size is modest, but it suggested the clinical utility of running a larger-scale MedCalc label recomputation in Phase 2.
3.2 Phase 2
Coverage and abstentions.
Using the supermajority-based labeling approach described in §2.4 yielded high-confidence labels for 887 out of 1,047 MedCalc test instances (85%) and for 4,593 out of 5,183 sampled train instances (89%). In these high-confidence labels, the abstention rate, i.e. how often the finalized label is $\mathsf{N/A}$, was 7.4% on the test set and 5.7% on the train set. We retain such $\mathsf{N/A}$ instances in the relabeled benchmark, as they correspond to clinically relevant “cannot be computed” cases rather than noise.
Estimated rate of likely mislabels.
Among the 887 test instances with high-confidence $\hat{y}^{\mathrm{new}}$, we find that 220 (24.8%) exceed the triage threshold (rel.err) defined in §2.4, plus another 66 (7.4%) are flagged as $\mathsf{N/A}$. These implies a total of 27.3% (on the 1,047) of the original test instances as likely mislabeled. This is still likely an underestimate, as rel.err is a conservative, symmetric lower bound on the underlying relative error with respect to the latent $y^{\star}$. Moreover, the 160 test instances for which the labeling agents could not reach a supermajority (and are therefore deferred) are plausibly hard cases where both $\hat{y}^{\mathrm{original}}$ and $\hat{y}^{\mathrm{new}}$ are unreliable, potentially pushing the true mislabel rate above 27.3%.
Cross-phase consistency check.
The close correspondence between Phase 1’s audit-based error rate (26.6%) and Phase 2’s recomputation-based estimate (27.3%) provides a sensitivity check. These phases differ meaningfully in their setup: Phase 1 audits labels with full metadata visibility, while Phase 2 recomputes without the anchoring. Still, they arrive at similar error estimates. While not constituting direct validation, this consistency suggests that the detected quality issues are robust to variations in prompting and workflow design, not dependent on a particular agentic configuration.
3.3 Phase 3
Table 1: Original MedCalc test labels ($\hat{y}^{\mathrm{original}}$) and our recomputed labels ($\hat{y}^{\mathrm{new}}$), compared against 50 physician-computed labels ($y^{\star}$). sMAPE, short for symmetric Mean Absolute Percentage Error, reports the average relative percentage deviation from physicians’ labels; it is computed on 34 / 33 instances where each $(\hat{y}^{\mathrm{original}},y^{\star})$ / $(\hat{y}^{\mathrm{new}},y^{\star})$ pair are both numerical. “Agreement” counts how many predicted labels in a group match the physician label up to a reasonable threshold. Square brackets are bootstrap $95\%$ confidence intervals.
Metric $(\hat{y}^{\mathrm{original}},y^{\star})$ $(\hat{y}^{\mathrm{new}},y^{\star})$
sMAPE (%) 72.7 [47.9, 99.1] 20.1 [8.0, 35.9]
Agreement 10/50 (20.0%) 37/50 (74.0%)
[…]